OpenStudy (anonymous):

f(x)=x^2+10x-25 find x intercept. Please help!

ganeshie8 (ganeshie8):

x intercepts occur when f(x) = 0

ganeshie8 (ganeshie8):

set the given function equal to 0 and solve x

ganeshie8 (ganeshie8):

f(x)= 0 x^2+10x-25 = 0

ganeshie8 (ganeshie8):

can you solve x ?

ganeshie8 (ganeshie8):

familiar with factoring a quadratic ?

OpenStudy (anonymous):

Not really im still learning factoring a quadratic

ganeshie8 (ganeshie8):

there is an easy trick

OpenStudy (anonymous):

I got (5,0)?

ganeshie8 (ganeshie8):

thats very close, but no

ganeshie8 (ganeshie8):

wait a sec , you might be correct if there is a typo in the original function

ganeshie8 (ganeshie8):

f(x) = x^2-10x+25

ganeshie8 (ganeshie8):

can you check + and - signs again ^^

OpenStudy (cj49):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\] a=1,b=10 and c=-25 plug in the values...can u do it from here...

OpenStudy (anonymous):

f(x)=-x^2+10x-25 find x intercept. The x^2= -x^2

ganeshie8 (ganeshie8):

haha! then you're right !! (5, 0) is the only xx intercept

OpenStudy (anonymous):

Thank you so much

ganeshie8 (ganeshie8):

-x^2+10x-25 = 0 x^2 - 10x + 25 = 0 (x-5)^2 = 0 x = 5

ganeshie8 (ganeshie8):

so the x intercept is (5, 0)

OpenStudy (anonymous):

And the y-intercept would be?

ganeshie8 (ganeshie8):

for y intercept, simply set x = 0 in the given function

OpenStudy (anonymous):

Do I set the problem up the same way?

ganeshie8 (ganeshie8):

f(x) = -x^2+10x-25 f(0) = ?

ganeshie8 (ganeshie8):

f(x) = -x^2+10x-25 f(0) = -0^2 + 10*0 - 25 = -25

ganeshie8 (ganeshie8):

so the y intercept is (0, -25)

OpenStudy (anonymous):

Im confused when it comes to adding square roots. will you always get the last number for y